This next question is reprinted from the textbook where it is question 17 at the end of the chapter. It describes an independent groups, two-factor study with two levels of A and two levels of B. Each cell has n = 10 particpants. Draw the 2x2 grid with Factor A (A1 and A2) on the left and Factor B (B1 and B2) across the top. The cell means are A1B1 = 3, A1B2 = 7, A2B1 = 5, and A2B2 is missing. You need to answer about the possible value of A2B2.

4a) What value for the missing mean would result in no main effect for Factor A?

4b) What value for the missing mean would result in no main effect for Factor B?

4c) What value for the missing mean would result in no interaction?

CAN YOU PLEASE EXPLAIN HOW YOU GOT THESE ANSWERS
Answer #1
B1 B2
A1 3 7
A2 5 a

let missing value is a:

4a)

for no main effect for Factor A : effect of A1 =effect of A2

A1B1+A1B2 =A2B1+A2B2

3+7 =5+a

a =missing value =5

4b)

for no main effect for Factor B : effect of B1 =effect of B2

A1B1+A2B1 =A1B2+A2B2

3+5 =7+a

a =missing value =1

4c)

for no interaction":

A1B1+A2B2 =A1B2+A2B1

3+a =5+7

a =missing value =9